OpenStudy (anonymous):

2x+4=-3x-11 What will the equation look like if I add 3x to both sides

OpenStudy (anonymous):

5x+4=-11

OpenStudy (anonymous):

2x+4=-3x-11 +3x +3x 5x+4=-11?

OpenStudy (anonymous):

5x + 4 = -11

OpenStudy (aravindg):

-11